40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 9 08:47:35 2012, steve, Update, SUS, ETMX damping restored 
    Reply  Mon Jan 9 13:19:09 2012, kiwamu, Update, SUS, ETMX damping restored 
       Reply  Thu Jan 12 23:19:56 2012, Koji, Update, SUS, c1iscex is fine now 
Message ID: 6194     Entry time: Thu Jan 12 23:19:56 2012     In reply to: 6181
Author: Koji 
Type: Update 
Category: SUS 
Subject: c1iscex is fine now 

c1iscex is working as before and the optic is damped.


What I checked

1. I went to the X-end rack. I found the io-chassis was turned off.

2. I shutdown c1iscex, turned off, and turned on everything. Again, we did not have any signal from the ADC into c1scx model.

However, I found that c1x01 indicates healthy ADC signals.
This means that the connection between the IOP and the c1scx model was wrong ==> Simulated Plant

3. Burtrestored X'mas eve snapshot. This restored the gains and matrices as well as C1:SCX-SIM_SWITCH
which switches the input between the real ADCs and simulated plant.

4. The signals came back to c1scx.

ELOG V3.1.3-